a block of iron mass 60kg must be heated from 20 degrees Celsius to 300 degrees Celsius. the specific heat capacity of iron is470J/kg k. the effeciencyof the system is 71%. the available supply voltage is. 230V. the process must be completed within 2 hours. if electricity from eskom is presentlyR1, 77 kilometers per hour. determine the value of current drawn from the source and the total cost of the process


1
Expert's answer
2021-06-08T22:19:02-0400

Part a

P=VI

"\frac{W}{t}=VI"

"\frac{mc \triangle T}{t}=VI"

"I=\frac{mc \triangle T}{Vt}=\frac{60 *470* 573-293 }{230 *2*60*60}=4.768 kW"


Part b

"1kWh = R1, 77"

"4.768kWh = \frac{R177*4.768}{1}=R843.96"